Consider three numbers x, y, and z that sum to 1 and also have their squares sum to 1. Find x, y, and z so that their product is a minimum.
Answer #1

given aty+z=1 y?+2? = 1 + 2 : the product is myz . 314 let frymez)=X+Y+ 2-1. . 2 g(ximiz) 2x²+y2+2 ²4 34 legrenge method. 218If t= -0 4 18 = 1/3 14=1 a 14 =18 pooduct Exyz = (hells) Cos) niin for minimum product t- tus) mar 19 = tz czando Thank you!